OpenStudy (anonymous):

Describe the end behavior of the rational function g(x) = - (6x+1)/(x+2)

OpenStudy (zzr0ck3r):

is this for calculus?

OpenStudy (anonymous):

Algebra 2

OpenStudy (anonymous):

With trig

OpenStudy (zzr0ck3r):

do you know how to find horizontal asymptotes?

OpenStudy (anonymous):

Doesn't it depend on the type of function it is??

OpenStudy (zzr0ck3r):

y = (-6x-1)/(x+2) solve for x y(x+2) = -6x-1 yx+2y = -6x-1 yx+6x=-1-2y x(y+6)=-(1+2y) x=-(1+2y)/(y+6)

OpenStudy (anonymous):

I thought the horizontal asymptote for this one would be 6

OpenStudy (zzr0ck3r):

the domain of this is the range of the starting function

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

this is the end behavior

OpenStudy (zzr0ck3r):

-6

OpenStudy (zzr0ck3r):

y=-6 is the end behaivior

OpenStudy (anonymous):

So the end behavior is the horizontal asymptote?

OpenStudy (anonymous):

or, the opposite of?

OpenStudy (zzr0ck3r):

as x approaches +-infinity, f(x) approaches -6

OpenStudy (agent0smith):

The highest powers determine end behaviour. Compare the highest powers of x in the numerator and denominator: \[\Large g(x) = \frac{ -6x-1 }{ x+2 }\] On the top, the highest power is the -6x. On the bottom, it's x. Then you divide them like so \[\large y= \frac{ -6x }{ x } = -6\]

OpenStudy (anonymous):

thank you guys